in linear equation in two variable
 
how to write an equation parallel to x axis passising through point (5,4)

Grade:10

1 Answers

Saurabh Kumar
9 years ago
ILine paralel to x-axis is (y – y1) = m(x-x1).
Since, slope of the line is ‘0’ as the line is paralle to x-axis.
So equation would be (y-4 = 0).

Good Luck..................................................................................................